I got the answer (I think) for this problem but how do I check it? I think the answer is angle A is 80 degrees and angle B is 66 degrees. Right or wrong?
Angles A and B are complimentary angles, and angle B is 14 degrees less than angle A. Find the measures of angle A and angle B

Answer by Nate(3500)

A + B = 90
A - 14 = B
------------------------------------
A + B - 90 = 0
A - B - 14 = 0
2A - 104 = 0
2A = 104
A = 52
--------------------
A + B = 90
52 + B = 90
B = 38
When they say that angles are complementary, the angles are 90 degrees.